Express 35 mm in cm, m and km
Express 35 mm in cm, m and km

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

We know that,
= 1 cm = 10 mm
Then,
= 1 mm = (1/10) cm
= 35 mm = (35/10) cm
= 3.5 cm
And,
= 1 meter = 100 cm
Then,
= 1 cm = (1/100) m
= 3.5 cm = (3.5/100) m
= (35/1000) m
= 0.035 m
Now,
= 1 km = 1000 m
Then,
= 1 m = (1/1000) km
= 0.035 m = (0.035/1000)
= 0. 000035 km

Thanks